OpenStudy (anonymous):

Solve the equation. 5x + 9 = 1 A.x = -3/5 B.x = 5/8 C.x = 8/5 D.x = -8/5

OpenStudy (anonymous):

well the easiest way would be to plug your options into the equation 5(-3/5) + 9 = ? 5(5/8) + 9 = ? 5(8/5) + 9 = ? 5(-8/5) +9 =? :D

OpenStudy (anonymous):

D -8/5 5x+9=1 5x=1-9 5x=-8 x=-8/5